# MOSS: Efficient and Accurate FP8 LLM Training with Microscaling and Automatic Scaling

### v1

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, including reasoning, language understanding, and generation (Achiam et al., 2023; Grattafiori et al., 2024; Liu et al., 2024; Adler et al., 2024). However, training these models, which often comprise billions of parameters, incurs substantial computational costs. By leveraging tensor quantization in deep neural networks, low-precision training has emerged as a promising solution to accelerate computation and reduce memory requirements. Among existing methods, BF16 training is currently the most widely adopted low-precision approach, supported by major large-scale training frameworks such as DeepSpeed (Rasley et al., 2020) and Megatron-LM (Shoeybi et al., 2019).
@@
However, despite the high fidelity, existing per-group FP8 quantization schemes apply scaling factors along the inner dimension (K) of GEMM executed on Tensor Cores, which incurs additional dequantization overhead on CUDA Cores. As illustrated in Figure 2, the per-group quantized COAT (Xi et al., 2024) significantly underperforms compared to the per-tensor approach used in Transformer Engine (Nvidia, 2025), due to significant dequantization and partial sum overhead within the main computation loop. Additionally, dynamically computing per-tensor scaling factors for weights introduces memory access and max-reduction overhead during training, diminishing the benefits of FP8 training.
@@
In this work, we present MOSS, a framework for efficient and accurate FP8 training that addresses the aforementioned limitations. For sensitive activations, MOSS introduces a progressive two-level quantization strategy: a global per-tensor scale in FP32 precision, combined with local microscaling using power-of-two factors. This design enhances numerical fidelity while reducing dequantization overhead in the GEMM kernel. For weights well-suited to per-tensor quantization, MOSS exploits the bounded update property of Adam-like optimizers (Kingma & Ba, 2014), where weight updates are inherently limited by the learning rate. Building on this, MOSS proposes an automatic scaling mechanism that predicts scaling factor evolution based on optimizer hyperparameters, thereby eliminating runtime max-reduction and its associated cost. Together, these innovations make MOSS a more efficient quantization framework for FP8 training, better suited to Tensor Cores while preserving training stability and accuracy.
We evaluate MOSS by training large language models, including OLMo-7B (Groeneveld et al., 2024) and LLaMA-2-7B (Grattafiori et al., 2024), under GPU resource-constrained settings, specifically, using only 8 Hopper GPUs for full-scale pretraining. Notably, although Hopper GPUs do not natively support microscaled data formats, we enable this functionality by implementing custom GEMM kernels using Triton OpenAI (2025). Experimental results demonstrate that MOSS achieves lossless accuracy on both pretraining and fine-tuning tasks, matching the performance of BF16 baselines. In terms of efficiency, MOSS delivers a $1.34\times$ end-to-end training speedup on OLMo-7B compared to BF16 and outperforms the state-of-the-art FP8 training framework, COAT, by $12.3\%$, significantly reducing the training cost of large-scale models. To further validate the efficiency of MOSS’s FP8 GEMM operations, we conduct an ablation study comparing its runtime against other FP8 GEMM kernels, including DeepGEMM from DeepSeek-V3 (Liu et al., 2024), which confirms the superior efficiency of MOSS’s GEMM kernel design.
@@ 2.1 FP8 and Microscaling Data Format
In FP8, each tensor $\bm{X}$ is mapped to its lower-precision counterpart $Q_{\bm{X}}$ via $Q_{\bm{X}}=\lceil\bm{X}/s\rfloor$, where $s$ is the FP32 scaling factor, enforcing all values to fall within the dynamic range of the FP8 format. This often necessitates the use of the less precise E5M2 format for certain tensors, such as gradients and activations, to accommodate wider ranges Nvidia (2025). The microscaling (MX) data format, initially proposed in (Darvish Rouhani et al., 2023), has been standardized by the Open Compute Project (Bita et al., 2023) with support from Microsoft, Intel, NVIDIA, and others. The primary difference between standard FP8 and MXFP8 quantization lies in the granularity of scaling. In particular, MXFP8 assigns a distinct scaling factor to each block of 32 consecutive values, enabling all values to be quantized using the higher-precision E4M3 format. A second key difference is in the representation of scaling factors: FP8 stores them in full-precision FP32, whereas MXFP8 uses an 8-bit power-of-two format (E8M0), significantly reducing storage and computational overhead.
@@ 2.2 Weight Update Rule
Optimizers play a crucial role in deep learning by updating model weights during training. The most common gradient-based optimizer is Adam/AdamW ( (Kingma & Ba, 2014; Loshchilov & Hutter, 2017)), which uses first-order $m$ and second-order momentum $v$ to achieve better convergence. Define $\nabla{\mathbf{W}}\mathcal{L}(\mathbf{W}_{t})$ as $\mathbf{g}_{t}$, the update rule for $l$-th layer at time step t using AdamW can be formulated as:
@@ 3.1 Two-Level MicroScaling
Overview. Per-group activation quantization in current FP8 training frameworks (Xi et al., 2024) introduces a critical drawback: the insertion of per-group scaling factors along the inner dimension K of GEMM operations incurs significant dequantization overhead. This overhead is executed on slower CUDA Cores rather than high-throughput Tensor Cores (see GEMM kernel illustration in Figure 3a). As a result, the dequantization steps within the main loop of FP8 GEMM become a major efficiency bottleneck. Specifically, on modern data center GPUs, the peak throughput of FP32 CUDA Cores is only 1.6% of that of FP8 Tensor Cores. As a result, dequantizing even a single partial sum in a per-group quantized FP8 GEMM can cost the equivalent of nearly 60 Tensor Core multiply-accumulate (MAC) operations. This imbalance causes the GEMM main loop to be dominated by slow dequantization on CUDA Cores, effectively nullifying the performance advantages of FP8 training.
@@
where $s_{i}$ is the scaling factor of subpartition $i$, $\bm{X}_{i}$ is the $i$-th group vector and $\Delta^{\text{FP8}}_{\max}$ is the maximum representable value in the chosen FP8 format. For common FP8 encodings (Micikevicius et al., 2023), $\Delta^{\text{E4M3}}_{\max}=448$ and $\Delta^{\text{E5M2}}_{\max}=57344$. In the second stage, we further quantize the per-group scale factors $s_{i}$ by separating them into E8M0 microscaling components $ss_{i}$ and a floating-point global component $s$,
@@
Following the MX format specification (Bita et al., 2023), we set the level-2 group size to $k_{2}=32$. To support low-latency access within the inner loop of Tensor Core MMAs, the level-2 scaling factors $ss_{i}$ are stored in a contiguous memory layout aligned with the expected access patterns. This organization ensures efficient memory utilization and prevents stalls during the GEMM main loop. For model weights, we apply per-tensor quantization with FP32 scaling factors to fully exploit Tensor Core performance. The quantized matrix multiplication in Figure 3b is performed as: $Q_{y}=Q_{w}\times(Q_{x}\ast ss_{x})$, where $Q_{x}$ is the FP8-quantized activation, scaled by E8M0-encoded level-2 factors $ss_{x}$, and $Q_{w}$ is the FP8-quantized weight tensor.
@@ 3.2 Automatic Scaling
Various scaling techniques have been proposed to adapt tensors to the limited dynamic range of the FP8 format (Nvidia, 2025; Blake et al., 2023). However, these methods often introduce training overhead due to additional memory accesses and max-reduction operations. For example, just-in-time scaling requires reading all FP32 values from High Bandwidth Memory (HBM) to compute the maximum absolute value for quantization. The resulting FP8 values are then written back to HBM, only to be read again during subsequent MMAs, incurring significant latency and bandwidth costs.
To address this inefficiency, we propose a novel automatic scaling approach that estimates the evolution of scaling factors for parameters throughout training. LLMs are typically trained using Adam or its variants (Kingma & Ba, 2014; Loshchilov & Hutter, 2017), which maintain additional copies of model weights, gradients, and first- and second-order momentums for parameter updates. A key advantage of these optimizers is their ability to produce stable updates, with changes to model parameters bounded by the step size $\eta$, regardless of gradient volatility. By exploiting this bounded-update property, we propose a novel automatic scaling scheme that can predict and adjust scaling factors ahead of time, eliminating the need for repeated memory reads and writes, and thus avoiding runtime overhead.
@@
To analyze the bound of $|\Delta_{t}|$, we first examine the term $|\mathbf{m}_{t}/\sqrt{\mathbf{v}_{t}}|$. During LLM training, both the first-order moment $\mathbf{m}_{t}$ and the second-order moment $\mathbf{v}_{t}$ are initialized to zero, i.e., $\mathbf{m}_{0}=\mathbf{0}$ and $\mathbf{v}_{0}=\mathbf{0}$. As training progresses, $\mathbf{m}_{t}$ becomes the exponential moving average (EMA) of the gradients, i.e., $\mathbf{m}_{t}=\mathbb{E}[\mathbf{g}]_{\text{EMA}}$, and $\mathbf{v}_{t}$ becomes the EMA of the squared gradients, i.e., $\mathbf{v}_{t}=\mathbb{E}[\mathbf{g}^{2}]_{\text{EMA}}$. Applying Jensen’s inequality (Abramovich et al., 2004) to the convex function $f(x)=x^{2}$ yields the relation $(\mathbf{m}_{t})^{2}\leq\mathbf{v}_{t}$, which in turn implies that $|\mathbf{m}_{t}/\sqrt{\mathbf{v}_{t}}|\leq 1$.
@@
In more typical cases, the effective update $|\Delta_{t}|$ is strictly smaller than the step size $\eta$. For example, when $1-\beta_{1}^{t}=\sqrt{1-\beta_{2}^{t}}$, and given that $|\hat{\mathbf{m}}_{t}/\sqrt{\hat{\mathbf{v}}_{t}}|\leq 1$, it follows directly that $|\Delta_{t}|<\eta$. In practice, during LLM training, it is common to have $1-\beta_{1}^{t}<\sqrt{1-\beta_{2}^{t}}$ because LLMs typically adopt smaller values for $\beta_{2}$ (e.g., $\beta_{2}=0.95$(Grattafiori et al., 2024; Groeneveld et al., 2024). This choice keeps the second-order moment estimate more responsive to recent gradients, helping to prevent both exploding updates and excessively slow learning. Overall, the effective parameter update at each timestep is approximately bounded by the step size $\eta$, i.e., $|\Delta_{t}|\leq\eta$. This behavior can be interpreted as implicitly defining a trust region around the current parameter values, within which the optimizer assumes the gradient estimate remains informative and reliable.
@@ 4.1 Experimental Setup
Setups. We train the language models using the AdamW optimizer (Loshchilov & Hutter, 2017), an extension of Adam (Kingma & Ba, 2014) with decoupled weight decay. Following common practice, we use decay rates of $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, and a weight decay of 0.1. The learning rate, beginning from $2e^{-4}$, follows a cosine decay schedule, with the final learning rate set to 10% of the peak value. Models are trained on a total of 22B tokens with a global batch size of 256, which is sufficient to preserve the Gaussian distribution of weight updates. The input sequence length is fixed at 2048 tokens, and a warm-up phase of 2,000 iterations is applied. Additional model configurations and training details are provided in the Appendix. Standard GPUs do not natively support the microscaling data format available on the newer Blackwell GPU series. To enable MXFP8 computation on standard GPUs, we implement custom matrix multiplication kernels using Triton (OpenAI, 2025). We use the proposed two-level microscaling method to quantize activations and per-tensor quantization with automatic scaling for model weights in linear layers.
Baselines and Datasets. We compare MOSS against BF16 training and the COAT baseline (Xi et al., 2024) to evaluate its effectiveness in both LLM pretraining and fine-tuning tasks. Transformer Engine (Nvidia, 2025) also adopts MXFP8 for quantization; however, its microscaling strategy is supported only on specific GPUs, which limits its broader applicability. For LLM pretraining, we train the OLMo-7B model on the Dolma corpus (Soldaini et al., 2024) and report perplexity on standard benchmarks including WikiText-103 (Merity et al., 2016), C4 (Raffel et al., 2020), and The Pile (Gao et al., 2020). For LLM fine-tuning, we focus on mathematical reasoning tasks by fine-tuning a LLaMA-2-7B model (Grattafiori et al., 2024) on the MAmmoTH dataset (Yue et al., 2023). We then evaluate the instruction-tuned model on three downstream benchmarks: Mathematics (Davies et al., 2021), GSM8K (Cobbe et al., 2021), and NumGLUE (Mishra et al., 2022).
4.2 LLM Pre-Training

### v2

@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities across diverse tasks, including reasoning, language understanding, and generation (achiam2023gpt; grattafiori2024llama; liu2024deepseek; adler2024nemotron). However, training these models, which often comprise billions of parameters, incurs substantial computational costs. By leveraging tensor quantization in deep neural networks, low-precision training has emerged as a promising solution to accelerate computation and reduce memory requirements. Among existing methods, BF16 training is currently the most widely adopted low-precision approach, supported by major large-scale training frameworks such as DeepSpeed (rasley2020deepspeed) and Megatron-LM (shoeybi2019megatron).
@@
However, despite the high fidelity, existing per-group FP8 quantization schemes apply scaling factors along the inner dimension (K) of GEMM executed on Tensor Cores, which incurs additional dequantization overhead on CUDA Cores. As illustrated in Figure 2, the per-group quantized COAT (xi2024coat) significantly underperforms compared to the per-tensor approach used in Transformer Engine (TE), due to significant dequantization and partial sum overhead within the main computation loop. Additionally, dynamically computing per-tensor scaling factors for weights introduces memory access and max-reduction overhead during training, diminishing the benefits of FP8 training.
@@
In this work, we present MOSS, a framework for efficient and accurate FP8 training that addresses the aforementioned limitations. For sensitive activations, MOSS introduces a progressive two-level quantization strategy: a global per-tensor scale in FP32 precision, combined with local microscaling using power-of-two factors. This design enhances numerical fidelity while reducing dequantization overhead in the GEMM kernel. For weights well-suited to per-tensor quantization, MOSS exploits the bounded update property of Adam-like optimizers (kingma2014adam), where weight updates are inherently limited by the learning rate. Building on this, MOSS proposes an automatic scaling mechanism that predicts scaling factor evolution based on optimizer hyperparameters, thereby eliminating runtime max-reduction and its associated cost. Together, these innovations make MOSS a more efficient quantization framework for FP8 training, better suited to Tensor Cores while preserving training stability and accuracy.
We evaluate MOSS by training large language models, including OLMo-7B (groeneveld2024olmo) and LLaMA-2-7B (grattafiori2024llama), under GPU resource-constrained settings, specifically, using only 8 Hopper GPUs for full-scale pretraining. Notably, although Hopper GPUs do not natively support microscaled data formats, we enable this functionality by implementing custom GEMM kernels using Triton triton. Experimental results demonstrate that MOSS achieves lossless accuracy on both pretraining and fine-tuning tasks, matching the performance of BF16 baselines. In terms of efficiency, MOSS delivers a $1.34\times$ end-to-end training speedup on OLMo-7B compared to BF16 and outperforms the state-of-the-art FP8 training framework, COAT, by $12.3\%$, significantly reducing the training cost of large-scale models. To further validate the efficiency of MOSS’s FP8 GEMM operations, we conduct an ablation study comparing its runtime against other FP8 GEMM kernels, including DeepGEMM from DeepSeek-V3 (liu2024deepseek), which confirms the superior efficiency of MOSS’s GEMM kernel design.
@@ 2.1 FP8 and Microscaling Data Format
In FP8, each tensor $\bm{X}$ is mapped to its lower-precision counterpart $Q_{\bm{X}}$ via $Q_{\bm{X}}=\lceil\bm{X}/s\rfloor$, where $s$ is the FP32 scaling factor, enforcing all values to fall within the dynamic range of the FP8 format. This often necessitates the use of the less precise E5M2 format for certain tensors, such as gradients and activations, to accommodate wider ranges TE. The microscaling (MX) data format, initially proposed in (darvish2023shared), has been standardized by the Open Compute Project (ocpmxfp8) with support from Microsoft, Intel, NVIDIA, and others. The primary difference between standard FP8 and MXFP8 quantization lies in the granularity of scaling. In particular, MXFP8 assigns a distinct scaling factor to each block of 32 consecutive values, enabling all values to be quantized using the higher-precision E4M3 format. A second key difference is in the representation of scaling factors: FP8 stores them in full-precision FP32, whereas MXFP8 uses an 8-bit power-of-two format (E8M0), significantly reducing storage and computational overhead.
@@ 2.2 Weight Update Rule
Optimizers play a crucial role in deep learning by updating model weights during training. The most common gradient-based optimizer is Adam/AdamW ( (kingma2014adam; loshchilov2017decoupled)), which uses first-order $m$ and second-order momentum $v$ to achieve better convergence. Define $\nabla{\mathbf{W}}\mathcal{L}(\mathbf{W}_{t})$ as $\mathbf{g}_{t}$, the update rule for $l$-th layer at time step t using AdamW can be formulated as:
@@ 3.1 Two-Level MicroScaling
Overview. Per-group activation quantization in current FP8 training frameworks (xi2024coat) introduces a critical drawback: the insertion of per-group scaling factors along the inner dimension K of GEMM operations incurs significant dequantization overhead. This overhead is executed on slower CUDA Cores rather than high-throughput Tensor Cores (see GEMM kernel illustration in Figure 3a). As a result, the dequantization steps within the main loop of FP8 GEMM become a major efficiency bottleneck. Similar dequantization-induced CUDA Core overhead has also been observed in inference systems such as QServe (lin2024qserve). Specifically, on modern data center GPUs such as the NVIDIA H100, the peak throughput of FP32 CUDA Cores is only 1.6% of that of FP8 Tensor Cores. As a result, dequantizing even a single partial sum in a per-group quantized FP8 GEMM can cost the equivalent of nearly 60 Tensor Core multiply-accumulate (MAC) operations. This imbalance causes the GEMM main loop to be dominated by slow dequantization on CUDA Cores, effectively nullifying the performance advantages of FP8 training.
@@
where $s_{i}$ is the scaling factor of subpartition $i$, $\bm{X}_{i}$ is the $i$-th group vector and $\Delta^{\text{FP8}}_{\max}$ is the maximum representable value in the chosen FP8 format. For common FP8 encodings (micikevicius2023ocp), $\Delta^{\text{E4M3}}_{\max}=448$ and $\Delta^{\text{E5M2}}_{\max}=57344$. In the second stage, we further quantize the per-group scale factors $s_{i}$ by separating them into E8M0 microscaling components $ss_{i}$ and a floating-point global component $s$,
@@
Following the MX format specification (ocpmxfp8), we set the level-2 group size to $k_{2}=32$. To support low-latency access within the inner loop of Tensor Core MMAs, the level-2 scaling factors $ss_{i}$ are stored in a contiguous memory layout aligned with the expected access patterns. This organization ensures efficient memory utilization and prevents stalls during the GEMM main loop. For model weights, we apply per-tensor quantization with FP32 scaling factors to fully exploit Tensor Core performance. The quantized matrix multiplication in Figure 3b is performed as: $Q_{y}=Q_{w}\times(Q_{x}\ast ss_{x})$, where $Q_{x}$ is the FP8-quantized activation, scaled by E8M0-encoded level-2 factors $ss_{x}$, and $Q_{w}$ is the FP8-quantized weight tensor.
@@ 3.2 Automatic Scaling
Various scaling techniques have been proposed to adapt tensors to the limited dynamic range of the FP8 format (TE; blake2023unit). However, these methods often introduce training overhead due to additional memory accesses and max-reduction operations. For example, just-in-time scaling requires reading all FP32 values from High Bandwidth Memory (HBM) to compute the maximum absolute value for quantization. The resulting FP8 values are then written back to HBM, only to be read again during subsequent MMAs, incurring significant latency and bandwidth costs.
To address this inefficiency, we propose a novel automatic scaling approach that estimates the evolution of scaling factors for parameters throughout training. LLMs are typically trained using Adam or its variants (kingma2014adam; loshchilov2017decoupled), which maintain additional copies of model weights, gradients, and first- and second-order momentums for parameter updates. A key advantage of these optimizers is their ability to produce stable updates, with changes to model parameters bounded by the step size $\eta$, regardless of gradient volatility. By exploiting this bounded-update property, we propose a novel automatic scaling scheme that can predict and adjust scaling factors ahead of time, eliminating the need for repeated memory reads and writes, and thus avoiding runtime overhead.
@@
To analyze the bound of $|\Delta_{t}|$, we first examine the term $|\mathbf{m}_{t}/\sqrt{\mathbf{v}_{t}}|$. During LLM training, both the first-order moment $\mathbf{m}_{t}$ and the second-order moment $\mathbf{v}_{t}$ are initialized to zero, i.e., $\mathbf{m}_{0}=\mathbf{0}$ and $\mathbf{v}_{0}=\mathbf{0}$. As training progresses, $\mathbf{m}_{t}$ becomes the exponential moving average (EMA) of the gradients, i.e., $\mathbf{m}_{t}=\mathbb{E}[\mathbf{g}]_{\text{EMA}}$, and $\mathbf{v}_{t}$ becomes the EMA of the squared gradients, i.e., $\mathbf{v}_{t}=\mathbb{E}[\mathbf{g}^{2}]_{\text{EMA}}$. Applying Jensen’s inequality (abramovich2004jensen) to the convex function $f(x)=x^{2}$ yields the relation $(\mathbf{m}_{t})^{2}\leq\mathbf{v}_{t}$, which in turn implies that $|\mathbf{m}_{t}/\sqrt{\mathbf{v}_{t}}|\leq 1$.
@@
In more typical cases, the effective update $|\Delta_{t}|$ is strictly smaller than the step size $\eta$. For example, when $1-\beta_{1}^{t}=\sqrt{1-\beta_{2}^{t}}$, and given that $|\hat{\mathbf{m}}_{t}/\sqrt{\hat{\mathbf{v}}_{t}}|\leq 1$, it follows directly that $|\Delta_{t}|<\eta$. In practice, during LLM training, it is common to have $1-\beta_{1}^{t}<\sqrt{1-\beta_{2}^{t}}$ because LLMs typically adopt smaller values for $\beta_{2}$ (e.g., $\beta_{2}=0.95$(grattafiori2024llama; groeneveld2024olmo). This choice keeps the second-order moment estimate more responsive to recent gradients, helping to prevent both exploding updates and excessively slow learning. Overall, the effective parameter update at each timestep is approximately bounded by the step size $\eta$, i.e., $|\Delta_{t}|\leq\eta$. This behavior can be interpreted as implicitly defining a trust region around the current parameter values, within which the optimizer assumes the gradient estimate remains informative and reliable.
@@ 4.1 Experimental Setup
Setups. We train the language models using the AdamW optimizer (loshchilov2017decoupled), an extension of Adam (kingma2014adam) with decoupled weight decay. Following common practice, we use decay rates of $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.95, and a weight decay of 0.1. The learning rate, beginning from $2e^{-4}$, follows a cosine decay schedule, with the final learning rate set to 10% of the peak value. Models are trained on a total of 22B tokens with a global batch size of 256, which is sufficient to preserve the Gaussian distribution of weight updates. The input sequence length is fixed at 2048 tokens, and a warm-up phase of 2,000 iterations is applied. Additional model configurations and training details are provided in the Appendix. Training is conducted on 8 NVIDIA H800 GPUs, which do not natively support the microscaling data format available on the newer Blackwell GPU series. To enable MXFP8 computation on standard GPUs such as NVIDIA Hopper, we implement custom matrix multiplication kernels using Triton (triton). We use the proposed two-level microscaling method to quantize activations and per-tensor quantization with automatic scaling for model weights in linear layers.
Baselines and Datasets. We compare MOSS against BF16 training and the COAT baseline (xi2024coat) to evaluate its effectiveness in both LLM pretraining and fine-tuning tasks. Transformer Engine (TE) also adopts MXFP8 for quantization; however, its microscaling strategy is supported only on NVIDIA Blackwell GPUs, which limits its broader applicability. For LLM pretraining, we train the OLMo-7B model on the Dolma corpus (soldaini2024dolma) and report perplexity on standard benchmarks including WikiText-103 (wiki103), C4 (C4), and The Pile (gao2020pile). For LLM fine-tuning, we focus on mathematical reasoning tasks by fine-tuning a LLaMA-2-7B model (grattafiori2024llama) on the MAmmoTH dataset (yue2023mammoth). We then evaluate the instruction-tuned model on three downstream benchmarks: Mathematics (mathematics), GSM8K (gsm8k), and NumGLUE (mishra2022numglue).
4.2 LLM Pre-Training
We begin by comparing the performance of OLMo-7B models (groeneveld2024olmo) pre-trained using FP8 mixed precision with those trained using BF16. The pre-training loss over training steps is shown in Figure 5. All models use identical training configurations and hyperparameters as specified in the official report. As illustrated in Figure 5, the loss curves of BF16, COAT, and our proposed method, MOSS, closely align, exhibiting nearly identical trends. These results clearly demonstrate that the proposed FP8 mixed-precision scheme can match the training performance of the widely adopted higher-precision BF16 baseline.
4.3 LLM Fine-Tuning
We further evaluate the proposed FP8 low-precision training scheme in the context of instruction tuning for LLMs. To ensure a fair comparison, we adopt the same fine-tuning setup as ToolBench (guo2024stabletoolbench), using the open-source LLaMA-2-7B (grattafiori2024llama) as the base model. Experiments are conducted on the math-focused MAmmoTH dataset (yue2023mammoth), fine-tuned for 5 epochs. During fine-tuning, MOSS matches the performance of the BF16 baseline, with loss curves that are either closely aligned or exhibit slightly improved stability (Figure 6). Downstream task accuracies are reported in Figure 6. These results further demonstrate the numerical precision and robustness of MOSS, confirming that the proposed FP8 quantization scheme generalizes well beyond pre-training and remains effective in fine-tuning settings.